OpenStudy (anonymous):

Given the functions f(x) = 4x2 − 1, g(x) = x2 − 8x + 5, and h(x) = –3x2 − 12x + 1, rank them from least to greatest based on their axis of symmetry. g(x), h(x), f(x) f(x), h(x), g(x) g(x), f(x), h(x) h(x), f(x), g(x)

OpenStudy (anonymous):

PLZZZZZZZZ HELPPPPP

OpenStudy (amoodarya):

\[f(x)=4x^2-1=4(x-0)^2-1 \rightarrow axis \space is \space x=0\\g(x)=x^2 − 8x + 5=(x-4)^2-16+5 \rightarrow axis \space is \space x=4\\h(x)= –3x^2 − 12x + 1=-3(x^2+4x)+1=-3((x+2)^2-4)+1 \rightarrow axis \space is \space x=-2\]

OpenStudy (amoodarya):

for the last one x=-2 is axis of symmetry